John is currently 4 times as old as his daughter, Linda. Six years from now, John will be 3 times as old as Linda. Calculate Linda's current age ?
Found 2 solutions by ikleyn, Edwin McCravy:
Answer by ikleyn(52776)

.
J = 4*L,            (1)

J + 6 = 3*(L+6).    (2)


Substitute (1) into (2). You will get

4*L + 6 = 3*(L+6)  ====>  4L +6 = 3L + 18  ====>  L = 18-6 = 12.


Answer.  Linda is 12 years old.  John is 4*12 = 48 years old.


There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.



Answer by Edwin McCravy(20054)

John is currently 4 times as old as his daughter, Linda.


J = 4•L

Six years from now, John will be 3 times as old as Linda.
(John + 6 = 3×(Linda + 6)



  J + 6 = 3•(L + 6)

So you have these two equations:

      J = 4•L
  J + 6 = 3•(L + 6)

Substitute 4•L from the first equation 
in place of J in the second equation:

4•L + 6 = 3•(L + 6)

You finish.

Edwin